# Generalizing fusion rules by shuffle: Symmetry-based classifications of nonlocal systems constructed from similarity transformations

### v2

@@ Abstract
Similarity transformation is a fundamental transformation, which enables one to obtain the Hermitian systems from pseudo-Hermitian systems by the invertible transformations[1, 2, 3]. This also implements a dual relationship between the nonunitary conformal field theories (CFTs) and the nonunitary CFTs describing the corresponding lattice models[4, 5]. One can see a similar (or the same) dual relationship in [6, 7, 8] related to the 3d3d3d-3d correspondence[9, 10, 11]. Related to the property of the effective vacuum or ground state, the Zamolodchikov $c$-theorem in unitary CFTs[12] has been generalized to $c_{\text{eff}}$-theorem[13] with a connection to the $PT$ symmetry in non-Hermitian systems[14, 15].
One can notice the nontrivial relationship between pseudo-Hermitian systems and the corresponding nonlocal Hermitian systems in the older studies on the realizations of chiral fermions[16, 17, 18, 19, 20, 21, 22]. The lattice realization of the chiral fermions is restricted by the Nielsen-Ninomiya theorem[23, 24], and the two possible ways to realize chiral fermions are breaking locality or Hermiticity. There also exists a lattice analogue in $1+1$-dimensions[25]. We note a recent review on chiral fermion[26] and those on non-Hermitian systems[27, 28, 29] as references.
@@
We also note the pioneering nonlocal model, Haldane-Shastry model[30, 31] and several recent research directions on non-local systems, numerical studies [32, 33, 34], long-range interacting integrable models [35, 36], extension of Lieb-Schultz-Mattis theorem[37, 38, 39, 40, 41, 42], and quantum information scrambling[43, 44, 45, 46, 47, 48, 49, 50, 51]. These studies have been developed in distinct research fields, with different techniques and languages. Hence, one may require some general method applicable to these models, such as symmetry analysis.
@@
In this paper, we construct and study a particular type of conserved charges, symmetry topological field theories (SymTFTs)[52], of the nonlocal Hermitian systems derived from pseudo-Hermitian systems by the similarity transformations. For readers interested in SymTFTs and generalized symmetries, we note several earlier works[53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65], recent reviews[66, 67, 68, 69] and two works[70, 71] by the authors and a collaborator. The fundamental algebraic structure is a generalization of the Verlinde formula [72, 73, 74], Eq. (14), and the corresponding SymTFTs and their modules are written as Eq. (20) and Eq. (28). Their implications to RGs are summarized in Fig. 1. However, the defect or boundary phenomena are more subtle because of the nonlocality, as we demonstrate in Sec. V.
@@ II Nonunitary CFTs and their symmetry
where $\{|\alpha\rangle\rangle\}$ is the Ishibashi states[84] satisfying the relation, $\widetilde{\langle\langle\alpha|}e^{\pi i\tau(L_{0}+\overline{L_{0}}-c/12)}|\beta\rangle\rangle=\chi_{\alpha}(\tau)\delta_{\alpha,\beta}.$ By applying the Verlinde formula, the symmetry operators act on the module in the following way[85],
@@
Moreover, the Cardy states satisfy the following relation[86, 87],
@@
As can be seen in Eq.(8), the vacuum, $I$, plays the central role. However, in nonunitary CFTs, there exists another fundamental object, the effective vacuum, $o$. Hence, one can naively expect the following to have some meaning,
@@
This matrix is less familar, but it has already appeared in [72, 73], and the section $8.19.$ of [74], and it has been known to be integers. Moreover, we demonstrate that there exists the corresponding algebra, $\mathbf{A}^{[o]}=\{\mathcal{Q}_{\alpha}^{[o]}\}$, which is ring isomorphic to $\mathbf{A}=\{\mathcal{Q}_{\alpha}\}$. Hence, if there exists a unitary CFT that has the same character as a nonunitary CFT, one can obtain the dual relation between their symmetries. Fortunately, there exists at least two series of orrespondence dual under the similarity, the correspondence between symplectic fermions and Dirac fermions[4] (or related $Sp(2n)_{k}\times Sp(2k)_{n}=SO(4nk)_{1}$ duality[88, 89]), and those between $M(2,2k+3)$ minimal models and $Osp(1,2)_{k}$ models (or related $W$-minimal models and $Osp(1,2n)_{k}$ models[8]). We expect the same type of dualities for the models with positive $\{S_{o,\alpha}\}_{\alpha}$, and we write the corresponding SymTFTs and their modules.
@@ III Characters and Symmetry operators by shuffle
$\chi_{\alpha}(\tau)=\langle\alpha,M|^{[o]}e^{2\pi i\tau(L_{0}^{[o]}-c^{[o]}/24)}|\alpha,M\rangle^{[o]}$ (17)
@@
Whereas we assumed the existence of the shuffle operation, a chiral and antichiral analogue of similarity transformations, but the combination $\eta^{[o]}\overline{\eta}^{[o]}$ can rigorously be identified as the similarity tranformation by applying the arguments in [1, 2, 3]. Corresponding to the above unitary representation, we also introduce the following projections,
@@
with $S_{o,\beta}\neq 0$. In many series of nonunitary CFTs, the condition $\{S_{o,\beta}>0\}_{\beta}$ is satisfied. Hence, the above symmetry operators are well-defined in many cases, and one can obtain the expression,
@@
It should be stressed that the condition $\{S_{j,\beta}\neq 0\}_{\beta}$ implies the invertibility of $\mathcal{Q}_{j}$ in linear algebra. It is well-known that the invertibility of the symmetry plays a significant role in studying gauging operations[64, 65], and our argument provides a generalization respecting the linear algebraic structure of quantum systems. Moreover, one can obtain the following relations,
@@ IV.1 Massless RG flow
First, we assume the massless RG or ring homomorphism $\rho$, $\rho:\mathbf{A}\rightarrow\mathbf{A^{\prime}}$[107, 118, 110]. In this setting, one can implement the ring homomorphism of the corresponding nonlocal unitary model as
@@
The unbroken symmetry $\mathbf{A}_{\text{ub}}$, such as $\mathbf{A}_{\text{ub}}\cap\text{Ker}\rho=\{0\}$ in the UV theory is transformed to $\iota(\mathbf{A}_{\text{ub}})$ and one can apply the analysis in [70, 71]. See FIG. 1 for intuition. For further studies, we introduce the following generalization of quantum dimensions (which can be called shuffled quantum dimensions),
@@ V Difficulty in boundary or defect phenomena
By choosing the left and right boundary conditions carefully for $Z_{\alpha,\beta}^{[o]}$, one can also avoid the negative amplitude. This nontrivial (or nonlocal) relation between the nonnegative amplitude condition and the boundary conditions can be interpreted as a signal of the nonlocality. For example, in the $M(2,5)$ model, one can obtain the same characters from the pairs $(Z_{o,o}^{[o]},Z_{I,o})$, $(Z_{o,I}^{[o]},Z_{I,I})$, but it is impossible to reproduce $Z_{o,o}$ from $Z^{[o]}_{\alpha,\beta}$, by taking the label $\alpha,\beta$ as that of the GW states. By introducting the unusual non-GW states $|o\rangle^{[o]}+|I\rangle^{[o]}$, $Z_{o,I+o}^{[o]}$ produces $Z_{o,o}$, but this breaks the correspondence between left and right boundaries in $Z_{o,o}$.
The phenomenon seems to be related to the boundary sensitity[128, 129, 130] or quantum skin effect[131] in non-Hermitian systems, and it is worth further study. We also note some works studying similar problems arising from the boundary conditions of pseudo-Hermitian systems[132, 133], and a pioneering work pointing out its relation to quantum groups[134].
@@ VI Conclusion and discussion
In this paper, we have studied the implication of the Galois shuffle operations in CFTs[90] and their implications as similarity transformations between pseudo-Hermitian systems and the corresponding nonlocal Hermitian systems. Essentially, this is a ring isomorphism between two theories, and the RGs of one side are compatible with the other side. However, when studying the implications for defects or boundaries, more careful treatment is necessary, and this difficulty is a natural consequence of the nonlocality.
@@ VII Acknowledgement
Appendix A Generalization and conjecture on similarity transformation

### v3

@@ Abstract
Similarity transformation, $\eta$, is a fundamental transformation, which enables one to obtain the Hermitian systems from pseudo-Hermitian systems by the invertible transformations [1, 2, 3]:
$H^{\prime}=\eta H\eta^{-1}.$ (1)
@@
where $H$ is the Hamiltonian of the original pseudo-Hermitian systems and $H^{\prime}$ is that of the corresponding Hermitian system [^1] In this manuscript, we assume that the spectrum of pseudo-Hermitian systems are real, but it will be possible to relax this assumption to the reality of partition functions.. This transformation also implements a dual relationship between the nonunitary conformal field theories (CFTs) and the nonunitary CFTs describing the corresponding lattice models [5, 6]. One can see a similar (or the same) dual relationship in [7, 8, 9] related to the 3d3d3d-3d correspondence [10, 11, 12]. It should be remarked that the similarity transformation itself is an intrinsically quantum operation, because it changes the role of vacuum in a CFT and the ground state (or effective vacuum) at the level of linear algebra[13, 14]. Because of this unusual aspect, it is difficult to represent the transformation in the standard Lagrangian formalism of quantum field theories (QFTs). Related to the property of the effective vacuum or ground state, the Zamolodchikov $c$-theorem in unitary CFTs [15] has been generalized to $c_{\text{eff}}$-theorem [16] with a connection to the $PT$ symmetry in non-Hermitian systems [17, 18].
@@
One can notice the nontrivial relationship between pseudo-Hermitian systems and the corresponding nonlocal Hermitian systems in the older studies on the realizations of chiral fermions [19, 20, 21, 22, 23, 24, 25]. The lattice realization of the chiral fermions is restricted by the Nielsen-Ninomiya theorem [26, 27], and the two possible ways to realize chiral fermions are breaking locality or Hermiticity. There also exists a lattice analogue in $1+1$-dimensions [28]. We note a recent review on chiral fermion [29] and those on non-Hermitian systems [30, 31, 32] as references. These research directions also imply the difficulty in describing the quantum models in the Lagrangian formalism of QFTs. We also note the pioneering nonlocal model, Haldane-Shastry model [33, 34] and several recent research directions on non-local systems, numerical studies [35, 36, 37], long-range interacting integrable models [38, 39], extension of Lieb-Schultz-Mattis theorem [40, 41, 42, 43, 44, 45], and quantum information scrambling [46, 47, 48, 49, 50, 51, 52, 53, 54]. These studies have been developed in distinct research fields, with different techniques and languages. Hence, one may require some general method applicable to these quantum mechanical models, such as symmetry analysis based on linear algebra.
On the other hand, there is a significant progress in the symmetry analysis in physics, known as generalized symmetry [55]. Roughly speaking, the symmetry in physics had been studied as a hierarchical structure of group, but this has been generalized to ring or algebra by treating the symmetry operators as conserved charges in quantum Hamiltonian systems[56]. In more recent technical terms, the symmetry has often been represented as a fusion category, but the identification between a fusion category and a (fusion) ring is more involved through topological Wick rotation (or modular $S$ transformation). Implicitly, this correspondence requires mapping between the quantum and classical systems. The underlying subtleties of this mapping have been clarified more recently in [57] in studying representations of the parafermionic lattice models [58]. Surprisingly, the parafermionic models represented by the parafermionic variables cannot be interpreted as two dimensional classical models, regardless of their apparent realization in the one dimensional quantum lattice models (equivalent to $Z_{N}$ generalization of Kitaev chain [59]). The resultant CFT can be interpreted as $Z_{N}$ extended models, and they are intrinsically quantum models. Moreover, the fusion category in literature usually does not permit noninteger coefficients before the anyonic objects, and this is intrinsically different from conserved charges of a quantum system, which should form a linear algebra.
This subtlty does not affect the analysis for a model when restricting our attention to a well-organized model with quantum/classical correspondence, such as a bosonic model described by a modular tensor category. However, it is not the case when studying their hierarchical structure and related extended models, which are intrinsically quantum. The relationship between the extended model and hierarchical structures has been proposed in [60] (One can see more detailed discussions and references in [61, 62]). From this point, the sandwich construction or symmetry-based representation of the correspondence between CFT and topological quantum field theory (TQFT) [63] should provide boundary states (expressed as linear vectors) in $D+1$ dimensional TQFTs and symmetry operators in the $D$ dimensional CFTs [64, 65, 62]. The resultant set of the conserved charges and their algebra have been known as symmetry topological field theories (SymTFTs) [66]. Related hierarchical structures have been studied as gapped domain wall [67, 68, 69, 70, 71] in TQFTs or massless renormalization group flow [72, 73, 74] in CFTs.
@@
In this paper, we construct and study a particular type of conserved charges, SymTFTs, of the nonlocal Hermitian systems derived from pseudo-Hermitian systems by the similarity transformations. For readers interested in SymTFTs and generalized symmetries, we note several earlier works [75, 76, 56, 77, 78, 79, 80, 81, 82, 60, 55, 83, 84], recent reviews [85, 86, 87, 88] and two works [89, 90] by the authors and a collaborator. The fundamental algebraic structure is a generalization of the Verlinde formula [13, 91, 92], Eq. (15), and the corresponding SymTFTs and their modules are written as Eq. (21) and Eq. (30). Their implications to RGs are summarized in Fig. 1. However, the defect or boundary phenomena are more subtle because of the nonlocality, as we demonstrate in Sec. V.
@@ II Nonunitary CFTs and their symmetry
where $\{|\alpha\rangle\rangle\}$ is the Ishibashi states [101] satisfying the relation, $\widetilde{\langle\langle\alpha|}e^{\pi i\tau(L_{0}+\overline{L_{0}}-c/12)}|\beta\rangle\rangle=\chi_{\alpha}(\tau)\delta_{\alpha,\beta}.$ We remind of the readers that the states and their linear duals are fundamental, and this is true for constructing the boundary states. By applying the Verlinde formula, the symmetry operators act on the module in the following way [102],
@@
In other words, the boundary states are defined as a natural module in linear algebra where the symmetry operators naturally act on. This aspect has been noticed in [102], and related phenomenologies have been developped in [97, 98, 99, 100, 61, 89, 90]. Moreover, the Cardy states satisfy the following relation [103, 104],
@@
As can be seen in Eq.(9), the vacuum, $I$, plays the central role. However, in nonunitary CFTs, there exists another fundamental object, the effective vacuum, $o$ where the conformal dimension $h_{o}$ is lowest in the theory. Hence, as an analog of local unitary RCFTs, one can naively expect the following to have some meaning,
@@
Indeed, as we see later, these coefficients appear as the coefficient of fusion algebra of symmetry operators or defects, whereas the defect realization requires more careful treatment. This matrix is less familiar, but it has already appeared in [13, 91], and the section $8.19.$ of [92], and it has been known to be integers. Eq. (15) has been studied in some categorical methods, but its realization and interpretation are less clear, because this can be outside of standard NIM-rep.
We demonstrate that there exists the corresponding algebra, $\mathbf{A}^{[o]}=\{\mathcal{Q}_{\alpha}^{[o]}\}$, which is ring isomorphic to $\mathbf{A}=\{\mathcal{Q}_{\alpha}\}$. Moreover, the new algebra $\mathbf{A}^{[o]}$ corresponds to the SymTFT of a nonlocal unitary CFT constructed from a local nonunitary CFT through the similarity transformation. Our primary method is linear algebra following the approach in [56]. Because of the fundamental theoretical strength of linear alegbra to quantum systems, our method applies to more general pseudo-Hermitian systems in principle. Hence, if there exists a unitary CFT that has the same character as a nonunitary CFT, one can obtain the dual relation between their symmetries. Fortunately, there exists at least two series of orrespondence dual under the similarity, the correspondence between symplectic fermions and Dirac fermions [105, 5] (or related $Sp(2n)_{k}\times Sp(2k)_{n}=SO(4nk)_{1}$ duality [106, 107]), and those between $M(2,2k+3)$ minimal models and $Osp(1,2)_{k}$ models (or related $W$-minimal models and $Osp(1,2n)_{k}$ models [8, 9]). We expect the same type of dualities for the models with positive $\{S_{o,\alpha}\}_{\alpha}$, and we write the corresponding SymTFTs and their modules. To our knowledge, the significance of the condition $\{S_{o,\alpha}>0\}_{\alpha}$ in nonunitary CFTs has never captured sufficient attention in the community[14, 90], but this condition is fundamental in constructing $\mathbf{A}^{[o]}$.
@@ III Characters and Symmetry operators by shuffle
$\chi_{\alpha}(\tau)=\sum_{M}\langle\alpha,M|^{[o]}e^{2\pi i\tau(L_{0}^{[o]}-c^{[o]}/24)}|\alpha,M\rangle^{[o]}$ (18)
@@
Whereas we assumed the existence of the shuffle operation, a chiral and antichiral analogue of similarity transformations, but the combination $\eta^{[o]}\overline{\eta}^{[o]}$ can rigorously be identified as the similarity tranformation by applying the arguments in [1, 2, 3]. At the level of linear algebra, the existence of similarity transformations is obvious, but their analytical properties are not studied except for the simplectic fermion [5]. Corresponding to the above unitary representation, we also introduce the following projections,
@@
with $S_{o,\beta}\neq 0$. In many series of nonunitary CFTs, the condition $\{S_{o,\beta}>0\}_{\beta}$ is satisfied[14]. Hence, the above symmetry operators are well-defined in many cases, and one can obtain the expression,
@@
It should be stressed that the condition $\{S_{j,\beta}\neq 0\}_{\beta}$ implies the invertibility of $\mathcal{Q}_{j}$ in linear algebra (not in category theory). It is well-known that the invertibility of the symmetry plays a significant role in studying gauging operations [83, 84] even in category theory, and our argument provides a generalization respecting the linear algebraic structure of quantum systems. Moreover, one can obtain the following relations,
@@ IV.1 Massless RG flow
First, we assume the massless RG or ring homomorphism $\rho$, $\rho:\mathbf{A}\rightarrow\mathbf{A^{\prime}}$ [68, 130, 71]. The ring homomorphism is a linear map satisfying the following relations,
$\rho(\alpha\times\beta)=\rho(\alpha)\times\rho(\beta)$ (27)
This relation implies that the fusion operation is deformed smoothly from one theory to another, and these smooth transtion properties have a close connection to the appearance of pseudo-Nambu-Goldstone (NG) modes (or Higgs modes) which should be gapped out gradually under the phase transitions [131, 132, 133, 134] (see also reviews [135, 136, 137]). In other words, at the level of linear algebra, the ring homomorphism determines the consequence after gapping out pseudo-NG modes, and the kernel of the map, $\text{Ker}(\rho)$, determines the pseudo-NG modes [89, 90, 138]. We note that the related correspondence between the massless RG and the level-rank duality [139, 140, 141, 142, 107] or coset construction [143, 144] has been studied in [145, 146, 147]. Related discussions can be seen in some recent references [148, 149]. It is worth noting that the simple object, $\alpha$, in the UV theory can be transformed to a nonsimple object $\sum_{\alpha^{\prime}}A^{\alpha^{\prime}}_{\alpha}\alpha^{\prime}$ in the IR theory, where $A$ is the coefficient. Because we have only assumed a linear algebraic structure of SymTFT, the coefficient can be noninteger [71, 89], and this is a natural generalization of anyon condensation in [150, 151]. In other words, if one does not permit the noninteger coefficient, the anyon condensation contains difficulties in classifying the existing quantum Hall hierarchy observed in experiments [152, 153] realizing the anomaly inflow mechanism[154, 155, 156, 157]. Moreover, this linear algebraic representation implies that the nonsimple object in UV theory can be simple in IR, and one can see the corresponding phenomena in the RG flow $\{SU(N)_{1}\}^{k}\rightarrow SU(N)_{k}$, because the abelian SymTFT, $\text{Rep}(\{SU(N)_{1}\}^{k})=\{Z_{N}\}^{k}$, flows to the nonabelian one [158, 159] (see [160, 161, 162, 163], for other examples in fractional quantum Hall hierarchies). Related phenomena have been studied in [164, 127, 165, 166, 167, 61, 168] with a connection to the Haldane conjecture [169, 170, 171] or anomaly matching. In other words, the emergence of nonabelian anyons or noninvertible symmetries can be understantood reasonable when assuming their underlying structure of coset constructions or level-rank dualities. Hence, nonsimple objects can be stabilized and simple object can split under the phase transition, and invertibility of objects can change dramatically.
In this setting, based on the above linear algebraic understanding, one can implement the ring homomorphism of the corresponding nonlocal unitary model as
@@
The unbroken symmetry $\mathbf{A}_{\text{ub}}$, such as $\mathbf{A}_{\text{ub}}\cap\text{Ker}\rho=\{0\}$ in the UV theory is transformed to $\iota(\mathbf{A}_{\text{ub}})$ and one can apply the analysis in [89, 90]. See FIG. 1 for intuition. The theoretical definition of the unbroken symmetry is straightforward [89], but its detection in particular massless flows is still in development. We note several recent works [172, 173, 174, 175]. However, we also point out that, whereas the detection of the unbroken symmetry in massless flows is important, this information cannot fix the corresponding homomorphism uniquely [89] without some other theoretical assumptions. For readers interested in the concrete descreptions of related phenomena, we note literature studying related domain wall or junction problems [176, 177, 178, 179, 180, 181, 182, 183, 184].
For further studies, we introduce the following generalization of quantum dimensions (which can be called shuffled quantum dimensions),
@@ IV.2 Smeared BCFT and massive RG flow
In a local theory, the GW states can be identified as the local boundary conditions generated by the Cardy boundary condition with some additional local degrees of freedom [185]. This is a consequence of the locality of boundary conditions and defects which are generated from the fusion of simple objects $a\times b\times c$.... The GW states provide a well-organized set of boundary conditions and the corresponding amplitudes in a local theory. But this is not the case for the nonlocal theory, and we discuss the problem in Sec. V. In other words, the insertion of symmetry operators and that of defects are intrinsically different in nonlocal systems, because the latter is constrained by the positive amplitude condition. The distinction between the symmetry operators and defects, called emanant symmetry, has been studied in [186, 187, 188, 189, 190], but our case is intrinsically different because the large volume limit does not resolve the discrepancies.
@@ V Difficulty in boundary or defect phenomena
By choosing the left and right boundary conditions carefully for $Z_{\alpha,\beta}^{[o]}$, one can also avoid the negative amplitude. This nontrivial (or nonlocal) relation between the nonnegative amplitude condition and the boundary conditions can be interpreted as a signal of the nonlocality. For example, in the $M(2,5)$ model, one can obtain the same characters from the pairs $(Z_{o,o}^{[o]},Z_{I,o})$, $(Z_{o,I}^{[o]},Z_{I,I})$, but it is impossible to reproduce $Z_{o,o}$ from $Z^{[o]}_{\alpha,\beta}$, by taking the label $\alpha,\beta$ as that of the GW states. By introducting the unusual non-GW states $|o\rangle^{[o]}+|I\rangle^{[o]}$, $Z_{o,I+o}^{[o]}$ produces $Z_{o,o}$, but this breaks the correspondence between left and right boundaries in $Z_{o,o}$. In other words, by constructing a nonlocal Hermitian system from a local pseudo-Hermitian system, there can exist unconventional boundary conditions outside of the GW states in nonlocal unitary BCFT. Inversely, there exists some cases where a nonlocal system with non-GW states can be mapped to a pseudo-Hermitian system with GW states.
The phenomenon seems to be related to the boundary sensitity [199, 200, 201] or quantum skin effect [202] in non-Hermitian systems, and it is worth further study. We also note some works studying similar problems arising from the boundary conditions of pseudo-Hermitian systems [203, 204, 205], and a pioneering work pointing out its relation to quantum groups [206].
@@ VI Conclusion and discussion
In this paper, we have studied the implication of the Galois shuffle operations in CFTs [14] and their implications as similarity transformations between pseudo-Hermitian systems and the corresponding nonlocal Hermitian systems. Essentially, this is a ring isomorphism between two theories, and the RGs of one side are compatible with the other side. However, when studying the implications for defects or boundaries, more careful treatment is necessary, and this difficulty is a natural consequence of the nonlocality. We remind that our discussion is based on linear algebra fundamental in studying quantum mechanical systems, and this aspect implies the necessity studying $\mathbb{C}$-linear category compatible with linear algebra.
@@ VII Acknowledgement
Appendix A $M(2,7)$ minimal models and $Osp(1,2)_{2}$ WZW model
In this subsection, we summarize the fusion coefficients of the $M(2,7)$ minimal model $N$ and those of the $Osp(1,2)_{2}$ WZW model, $N^{[o]}$. The $M(2,7)$ minimal model has three primary fields, $\{I,\phi,o\}$ with the conformal dimensions $\{h_{I}=0,h_{\phi}=-2/7,h_{o}=-3/7\}$. The conformal dimension and the effective conformal dimension are c=68/7c=-68/7 and $c_{\text{eff}}=4/7$. Hence, by considering the effective central charge $c_{\text{eff}}$ and effective conformal dimensions $h^{[o]}_{\alpha}$, one can check its correspondence to $Osp(1,2)_{2}$ WZW model. The modular $S$-matrix for this model is,
$S=\frac{2}{\sqrt{7}}\begin{pmatrix}\sin\frac{2\pi}{7}&-\sin\frac{4\pi}{7}&\sin\frac{6\pi}{7}\\ -\sin\frac{4\pi}{7}&\sin\frac{8\pi}{7}&-\sin\frac{12\pi}{7}\\ \sin\frac{6\pi}{7}&-\sin\frac{12\pi}{7}&\sin\frac{18\pi}{7}\end{pmatrix}$. (34)
One can construct $\{\mathcal{Q}_{\alpha}\}$ and $\{\mathcal{Q}_{\alpha}^{[o]}\}$ from these modular data. We summarize the corresponding fusion coefficients in the following table.
$\alpha$ $\beta$ $\gamma$ ${N^{\gamma,[I]}}_{\alpha\beta}$ ${N^{\gamma,[o]}}_{\alpha\beta}$
II II II 11 0
II II $\phi$ 0 1-1
II II $o$ 0 11
II $\phi$ II 0 1-1
II $\phi$ $\phi$ 11 11
II $\phi$ $o$ 0 0
II $o$ II 0 11
II $o$ $\phi$ 0 0
II $o$ $o$ 11 0
$\phi$ II II 0 1-1
$\phi$ II $\phi$ 11 11
$\phi$ II $o$ 0 0
$\phi$ $\phi$ II 11 11
$\phi$ $\phi$ $\phi$ 0 1-1
$\phi$ $\phi$ $o$ 11 11
$\phi$ $o$ II 0 0
$\phi$ $o$ $\phi$ 11 11
$\phi$ $o$ $o$ 11 0
$o$ II II 0 11
$o$ II $\phi$ 0 0
$o$ II $o$ 11 0
$o$ $\phi$ 0 0 0
$o$ $\phi$ $\phi$ 11 11
$o$ $\phi$ $o$ 11 0
$o$ $o$ II 11 0
$o$ $o$ $\phi$ 11 0
$o$ $o$ $o$ 11 11
More generally, one can observe the matching of the effective central charge $c_{\text{eff}}=2k/(2k+3)$ of $M(2,2k+3)$ minimal model and the central charge $c$ of the $Osp(1,2)_{k}$ WZW model. Phenomenologically, the $M(2,2k+3)$ minimal model corresponds to the bosonic-local-nonunitary CFT, and the $Osp(1,2)_{k}$ WZW model corresponds to the fermionic-nonlocal-unitary CFT. By interpreting the $M(2,2k+3)$ minimal model as (multicritical) Lee-Yang model [210] and the $Osp(1,2)_{k}$ WZW model as a supersymmetric model, this is analogous to dimensional reduction (or its breaking) [211]. The representation theories and fusion constants of the $Osp(1,2)_{k}$ WZW model are discussed in [212, 213]. We also note that the correspondence between the fermionic and bosonic representations of the characters is known as the Andrews-Gordon type identity [214, 215], a particular type of the Rogers-Ramanujan identities (see also [216, 217, 218] and literature therein). More recently, related phenomena in Chern-Simons theories have been studied in [142, 107]. Related phenomenologies have been summarized in [219] by the first author, and we only note a few earlier works [220, 221, 222, 223] here.
Appendix B Generalization and conjecture on similarity transformation